can u show me a way to do this fraction 3
_ v = 6
4

Answer by mananth(16946)

3/4 *v =6
multiply both sides of the equation by 4 to eliminate the denominator
3/4 * 4 *v= 6*4
3*v= 24
Divide both sides of the equation by 3 to eliminate 3 so that only v remains on the left side of the equation
3v/3= 24/3
v= 8